OpenStudy (anonymous):

Find the slope of the line passing through the points (2, 5) and (0, 4).

OpenStudy (anonymous):

slope of the line joining two points (x1,y1) and (x2,y2) is given by m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

thus here reqd slope= (4-5)/(0-2)=-1/-2=1/2

OpenStudy (anonymous):

is the answe 9/2